# Critical Quantum Sensing: a tutorial on parameter estimation near quantum phase transitions

### v1

@@ VI.13 Finite temperature precision comparison of single and multiparameter cases
which enters the QCRB Eq. (20) for multiparameter estimation. The diagonal elements of this inverse matrix quantify the minimal achievable variance for each parameter, accounting for correlations between them. In contrast, for single-parameter estimation of our target $\omega$ the QCRB simplifies to $\text{Var}(\omega)\geq\left[\mathcal{I}_{\omega}\right]^{-1}$, the inverse variance. To fairly compare the two approaches, we must relate these quantities. This involves relating the single parameter QFI, $\mathcal{I}_{\omega}$, to the inverse of the corresponding multiparameter variance from the QCRB: $\left(\mathcal{I}_{gg}^{\beta}/\text{det}\left[\hat{\mathcal{I}}_{\text{LZ}}^{\beta}\right]\right)^{-1}$.
@@
These features can be understood on a more quantitive basis in panel Fig. 11(c), where the single and multiparameter precision is shown as a function of temperature in the black solid line and dash dotted orange line, respectively. Here $g=0.8$ and $\omega=1$. The limit $T\!\to\!0$ reveals the single-parameter sensitivity remains finite whereas the orange dashed line (representing multiparameter sensitivity) tends to zero. When $T\sim\Delta$ the thermal fluctuations enhance the single parameter QFI and restore precision in the multiparameter case; however the magnitude of the multiparameter case is lower-bounded by the black line.
@@ VI.14 Attainability of multiparameter quantum Cramér-Rao bound
Thus far, we have encountered two key subtleties specific to the multiparameter estimation scenario. First, the QFIM can become singular, indicating that the bare parameters cannot be simultaneously estimated due to underlying dependencies between them. Second, there exists an inherent trade-off between the single and multiparameter estimation strategies. In the multiparameter setting, the presence of additional unknown quantities—while potentially offering a more realistic characterization, can degrade the estimation precision of a target parameter (in our case the energy splitting $\omega$).
Another important issue arises in relation to the attainability of the QCRB. While in single-parameter estimation the bound provided by the QFI is always asymptotically saturable using optimal measurements, this is not generically true in the joint estimation of multiple parameters. This can arise when the SLD operators $\hat{L}_{i}$ for different parameters are incompatible: meaning their non-commutativity prevents simultaneous optimal measurement and thus induces a fundamental precision compromise that cannot be overcome asymptotically [115, 106, 109, 172, 173, 174, 175, 105, 108, 176, 150, 97, 177, 110, 95, 112]. We will unpack these ideas for the presented LZ model. The SLD operators $\hat{L}_{\omega}$ and $\hat{L}_{g}$ can be computed using Eq. (21) and take the concise analytical form:
@@
meaning there is no mutual optimal measurement basis to estimate both parameters simultaneously. This means that fundamentally quantum precision trade-off exist as information is encoded in non-orthogonal directions in Hilbert space which “interfere” with each other. This interference prohibits a single measurement (or single-copy probe) from saturating the QCRB [110, 178, 179, 112]. Importantly, non-commutativity does not always imply incompatibility. In the LZ model we find a particularly interesting condition holds:
@@ VII Quantum phase transitions, criticality, and critical metrology: many-body Ising model
The Landau–Zener model has served as a minimal framework for understanding the fundamental principles by which quantum critical systems acquire enhanced sensing capabilities. Having established these operating principles, we now turn to many-body quantum systems capable of supporting bona fide quantum phase transitions. For this purpose, we introduce the general form of a spin-chain Hamiltonian,
$\hat{H}=\omega\sum_{i=1}^{N}\hat{\sigma}_{z}^{(i)}-\sum_{i\neq j=1}^{N}g_{ij}\hat{\sigma}_{x}^{(i)}\hat{\sigma}_{x}^{(j)}\;,$ (123)
where $\omega$ is the transverse field (the parameter to be estimated) and $g_{ij}$ denotes the interaction strength between the $i$th and $j$th spin. While Eq. (123) encompasses a broad class of interacting spin models, in what follows we restrict our analysis to two limiting cases. By first considering uniform nearest-neighbour interactions, $g_{ij}=g$, we obtain the one-dimensional transverse field Ising model (TFIM):
@@
  • Paramagnetic phase ($|\omega|>|g|$): In this phase, spins align along the transverse field direction $\hat{z}$ with short-range correlations. Connected spin-spin correlations decay exponentially with distance, reflecting the absence of long-range order. Here, correlations between the spins are small, and the ground state is close to a product state. This is also known as the disordered phase.
  • Ferromagnetic phase ($\omega<g$): In this phase, spins tend to align along one of the two directions $\pm\hat{x}$, reflecting the $\mathbb{Z}_{2}$ spin-flip symmetry of the Hamiltonian. For a finite system, the true ground state remains symmetry-preserving and is given by a superposition of the two ferromagnetic configurations. In the thermodynamic limit, however, spontaneous symmetry breaking occurs, and the system selects one of the ordered states. Connected correlations $\langle\hat{\sigma}_{x}^{(i)}\hat{\sigma}_{x}^{(i+r)}\rangle-\langle\hat{\sigma}_{x}^{(i)}\rangle\langle\hat{\sigma}_{x}^{(i+r)}\rangle$ saturate to a finite value as $r\to\infty$, signalling long-range order. This phase is also known as the ordered phase and supports a degenerate groundstate manifold in the thermodynamic limit.
@@
The TFIM represents a class of system which is exactly solvable via a Jordan-Wigner transformation by mapping spins to spinless fermions. A detailed discussion of the exact solution can be found in the following useful references [180, 181, 182, 163, 183, 184]. After Fourier transforming to momentum space, the Hamiltonian decouples into independent momentum-space pairs (k,k)(k,-k). These can be viewed as elementary quasiparticle modes that are always created or annihilated in pairs of opposite momentum. Within each such sector, the relevant states are the $|0\rangle$ (no quasiparticle excitations) and $\hat{c}^{\dagger}_{k}\hat{c}_{-k}^{\dagger}|0\rangle$ (quasiparticle pair with opposite momenta). This effectively reduced the problem to a two-level system for every momentum-$k$ mode. To make this structure explicit, we introduce fermionic two-component spinors
@@
where in this language, the dynamics of each (k,k)(k,-k) block resembles a LZ type model whose pseudospin representation encodes the presence or absence of quasiparticle excitations. Here the allowed momenta are given by:
@@
The $1/N$ scaling of the gap is a hallmark of the critical point in a finite-size system, reflecting how the system size itself sets the dominant length scale when the correlation length diverges. In this true many-body system, the closing of the gap emerges collectively from the full set of momentum $k$ modes. In contrast to this, the LZ model provides a simplified analogue: its single avoided crossing mimics the role of these momentum-resolved level crossings, thereby capturing the essence of the true many-body critical behaviour. In this picture, each momentum block behaves like an effective two-level system whose groundstate can be written as
@@
$|\psi_{GS}\rangle=\bigotimes_{k>0}|\psi_{GS}\rangle_{k}\;,$ (132)
@@ VII.1 Criticality as a resource for sensing
We here seek to explore the underlying characteristics of critical systems which make them an appealing platform in the context of quantum sensing. Quantum criticality is driven by quantum fluctuations, as opposed to thermal fluctuations in classical phase transitions. At zero temperature some change in a Hamiltonian parameter, denoted $g$, causes a fundamental change in the groundstate properties of the system. Notably, at or near the close vicinity of the critical point $g\equiv g_{c}$, the quantum fluctuations begin to dominate. This is accompanied by non-analytical behaviour in the groundstate which is reflected in physical observables, and qualitative changes in the nature of groundstate correlations. At finite temperature, there is a delicate interplay between thermal and quantum fluctuations.
In such cases, there exists a finite “quantum critical” regime where quantum fluctuations still dominate before thermal effects become significant. This interplay has important consequences for the robustness and practicality of quantum critical sensors. Furthermore, quantum criticality and entanglement are closely connected [185, 186, 187, 188, 189], although no unified theory links them fully. At first glance, the genuine multipartite entanglement structure of groundstates in quantum critical systems makes them attractive from a metrological perspective. For example, we have previously encountered the utility of leveraging entangled states in the context of metrology through Ramsey interferometry using entangled GHZ states in Sec. IV. In that instance, entanglement was an explicit resource facilitating enhanced sensing capabilities. However, as we will motivate in the following, it is the confluence of the entanglement structure in conjunction with the non-analytical behaviour near quantum critical points that underlies their utility as quantum sensors.
@@
in the traverse field $\omega$, manifest in large deviations of the groundstate $|\psi_{GS}(\omega)\rangle$, with $\delta\omega$ representing a small perturbation. For this purpose, we will consider the fidelity of pure states: a geometric object quantifying the overlap between two quantum states $|\psi_{GS}(\omega)\rangle$ and $|\psi_{GS}(\omega^{\prime})\rangle$ defined as [190]
@@
The fidelity measures the closeness between the two quantum states. If the state is unchanged then $F(\omega,\omega^{\prime})=1$ whereas orthogonal states are given as $F(\omega,\omega^{\prime})=0$. Contextually, the fidelity serves as a useful diagnostic tool for quantifying the variations $\delta\omega$ in the groundstate properties across the transition point [26, 191, 24, 192, 193, 25, 194, 195].
Using the exact solution of the TFIM the groundstate fidelity takes the concise form [196]
@@
for arbitrary system size $N$, where the product runs over momentum blocks of the Hamiltonian $\hat{H}_{k}$. Figure 12(a) shows the groundstate fidelity (135) landscape as a function of detuning away from the critical point, $\delta g_{c}=g-g_{c}$, and system size $N$, for fixed $\omega=1$ and $\delta\omega=0.01$. The colour scale highlights two qualitatively different regimes: the extended yellow regions correspond to fidelity values very close to unity, indicating that the groundstate remains largely unchanged under the small parameter shift. In contrast, the black/purple valley centred at the critical point, $\delta g_{c}=0$, marks a sharp loss in the fidelity (the groundstate changes markedly) in the quantum-critical region. As the system size $N$ increases, this valley deepens and narrows: reflecting a growing instability of the groundstate to perturbations at criticality. We further remark that in the thermodynamic limit, the fidelity can drop to zero irrespective of how small or large the deviation away $\delta\omega$ is. This phenomena is known as the Anderson orthogonality catastrophe [197].
@@ VII.2 Fidelity susceptibility
The connection between the dramatic loss of fidelity at the quantum critical point and indication these features are useful for metrology can be made rigorous however, by considering groundstate properties which are differentiable in parameter space. In particular, the fidelity between two infinitesimally close groundstates, for which $\delta\omega\ll 1$, can be expanded as [191, 26, 144, 194, 198, 196, 199]
@@
$\chi_{F}(\omega)$ $=-\frac{\partial^{2}F}{\partial(\delta\omega)^{2}}\biggr\rvert_{\delta\omega=0}\;,$ (137)
@@
The fidelity $F(\omega,\omega^{\prime})$ is a geometric measure of how different the groundstates are locally in parameter space, whereas the fidelity susceptibility $\chi_{F}(\omega)$ represents how sensitive the ground state is to infinitesimal variations in $\omega$. The fidelity susceptibility behaves like a response function in linear response theory. Concretely, it can be interpreted as a generalized susceptibility quantifying the ground state “response” to an infinitesimal change in a parameter $\omega$ in its state overlap, as opposed to its change in expectation value [200, 201, 192, 196, 26].
This is important as in quantum critical theory, linear response formalism is the standard tool to study how observables react to weak perturbations, with susceptibilities encoding correlations and critical scaling. In this sense, the fidelity susceptibility plays the role of a geometric analogue, quantifying the response of the ground state itself to infinitesimal parameter changes. In the TFIM the fidelity susceptibility can be computed directly
@@
We illustrate the behaviour of the fidelity susceptibility in Fig. 12(c), where it is shown as a function of the detuning from criticality, $\delta g_{c}=g-g_{c}$, and the system size $N$. In this analysis, the transverse field is fixed at $\omega=1$. The color-scale spans several orders of magnitude on a logarithmic scale: from dark purple (low susceptibility) of order $10^{-2}$ to bright yellow (high susceptibility) of order $10^{7}$ marking the large changes in the fidelity susceptibility. Several key features are immediately visible:
@@
  • Broadening away from criticality: For finite detuning ($|\delta g_{c}|>0$), the susceptibility grows much slower, indicating that the system is far less susceptibility to parameter variations away from criticality.
@@
Figure 12: Criticality as a metrological resource in the TFIM. (a) Ground state fidelity $F(\omega,\omega+\delta\omega)$ for $\delta\omega=0.01$ versus system size $N$ and detuning away from critical point $\delta g_{c}=g_{c}-1$, with $\omega=g_{c}=1$. The sharp dip at $\delta g_{c}=0$ reflects maximal state distinguishability. Note the log-scale on the $y$-axis. (b) Fidelity at criticality $g_{c}=\omega=1$ decaying rapidly with system size NN (log-scale), signalling the groundstate dramatically changing. (c) Fidelity susceptibility $\chi_{F}(\omega)$ (color in log-scale) as a function of system size NN (log-scale) and detuning from criticality $\delta g_{c}$. $\chi_{F}(\omega)$ diverges at the critical point marking the fact that the groundstate changes dramatically. (d) Scaling of the fidelity susceptibility $\chi_{F}(\omega)$ at the critical point $g_{c}=\omega=1$. It shows a power-law scaling $\chi_{F}(\omega=g_{c})\propto N^{\mu}$ illustrating that diverging nature at the critical point.
We can interrogate the divergence of the fidelity susceptibility at the critical point quantitatively by performing a scaling analysis. Setting $g=1$ for simplicity, we find that
$\chi_{F}$ $=\frac{1}{4}\sum_{k}\frac{\sin^{2}(k)}{\left[1-\cos(k)\right]^{2}}\;,$ (141)
$\eqsim\frac{N}{2}\int_{\pi/N}^{\pi(N-1)/N}\frac{1}{k^{2}}dk\;,$ (142)
@@
From this it is evident that $\chi_{F}(\omega)$ scales quadratically with system size $N$. Importantly, the fidelity susceptibility is not merely a heuristic measure of the groundstate sensitivity to infinitesimal variations: it can also be related to universal properties of a quantum critical point. Its scaling encodes universal properties of the system such as the correlation length exponent $\nu$ from $\xi\sim|g-g_{c}|^{-\nu}$ and spatial dimension $d$ when evaluated at the transition point. In particular it has been shown [205, 194, 203, 202] that close to a quantum critical point, the fidelity susceptibility follows a universal power-law behaviour,
@@
where the exponents depend only on the universality class of the transition, and not on the microscopic details of the Hamiltonian. The connection between the universal scaling functions of critical systems and the fidelity susceptibility is of direct consequence in metrology. This is because the fidelity susceptibility can in fact be seen as a specific component of the quantum geometric tensor, previously introduced in Sec VI.4. Utilizing this connections, one finds that
@@
This makes the rigorous connection between the dramatic changes groundstates undergo at quantum phase transitions, diverging susceptibilities, and ensuing metrological utility. Concretely, the fidelity susceptibility precisely captures the inherent change in groundstate properties that emerge near the quantum critical point. Furthermore, $\chi_{F}$ directly encodes universal scaling features of a given phase transition. In turn, through the QCRB, the QFI describes the fundamental attainable precision in parameter estimation, and is directly related to the fidelity susceptibility. This therefore suggests that optimal measurement strategies utilizing critical quantum probes inherently leverage the diverging nature of critical systems. From this rather simple illustrative system, we have demonstrated from a geometric perspective how features of critical systems naturally translate into Heisenberg (quadratic) like scaling from a metrological point of view. There is an intrinsic connection between divergent susceptibilities acting as a metrological amplifier in critical systems. Importantly, the presented arguments apply strictly at zero-temperature. In the general case, we expect there to be a delicate interplay between thermal and quantum fluctuations, and the resulting metrological utility of quantum critical sensors as has been seen in the case of LZ model in Sec. VI.6.
The coveted Heisenberg scaling to precision in quantum measurements can be attained in both the Ramsey interferometric scheme utilizing entangled GHZ states (as we have seen in Sec. IV) and in the TFIM at the quantum critical point. Nevertheless, the underlying physical principles facilitating the attainability of this precision are fundamentally different. To understand the distinction, we can view both the Ramsey scheme and the TFIM as representative examples of regular quantum metrology and critical quantum metrology, respectively. While the GHZ-based approach achieves enhanced sensitivity through maximally entangled probe sates and collective phase accumulation, critical quantum metrology explicitly leverages the diverging groundstate properties and long-range correlations near quantum critical points as the driving metrological resource. The appealing aspect of critical quantum metrology is that such schemes are, for the most part, basically unencumbered by the practical necessities required to prepare highly entangled states explicitly and furthermore, are more robust against decoherence effects which deteriorate any perceptible quantum advantage.

### v2

@@ VI.13 Comparison of precision at finite temperature for single and multiparameter cases
which enters the QCRB Eq. (20) for multiparameter estimation. The diagonal elements of this inverse matrix quantify the minimal achievable variance for each parameter, accounting for correlations between them. In contrast, for single-parameter estimation of our target $\omega$ the QCRB simplifies to $\text{Var}(\omega)\geq\left[\mathcal{I}_{\omega}\right]^{-1}$, the inverse of the variance. To fairly compare the two approaches, we must relate these quantities. This involves relating the single parameter QFI, $\mathcal{I}_{\omega}$, to the inverse of the corresponding multiparameter variance from the QCRB: $\left(\mathcal{I}_{gg}^{\beta}/\text{det}\left[\hat{\mathcal{I}}_{\text{LZ}}^{\beta}\right]\right)^{-1}$, which we denote as $\mathcal{I}_{\omega}^{\text{MP}}$ to be succinct.
@@
These features can be understood on a more quantitive basis in panel Fig. 11(c), where the single and multiparameter precision are shown as a function of temperature for $g=0.8$ and $\omega=1$. In the limit $T\!\to\!0$, the single-parameter sensitivity remains finite, whereas the multiparameter sensitivity vanishes. When $T\sim\Delta$, thermal fluctuations enhance the single-parameter QFI and restore precision in the multiparameter case. Notably, the multiparameter precision is always lower-bounded by the single-parameter value.
@@ VI.14 Attainability of multiparameter quantum Cramér-Rao bound
Thus far, we have encountered two key subtleties specific to the multiparameter estimation scenario. First, the QFIM can become singular, indicating that the bare parameters cannot be simultaneously estimated due to underlying dependencies between them. Second, there exists an inherent trade-off between the single and multiparameter estimation strategies. In the multiparameter setting, the presence of additional unknown quantities, while potentially offering a more realistic characterization, can degrade the estimation precision of a target parameter (in our case the energy splitting $\omega$).
Another important issue concerns the attainability of the QCRB. While in single-parameter estimation the bound provided by the QFI is always asymptotically saturable using optimal measurements, this is not generically true in the joint estimation of multiple parameters. When the SLD operators $\hat{L}_{i}$ corresponding to different parameters are incompatible—that is, they do not commute—simultaneous optimal measurement is impossible. This non-commutativity induces a fundamental trade-off in precision that cannot be overcome, even asymptotically. [127, 118, 121, 185, 186, 187, 188, 117, 120, 189, 163, 109, 190, 122, 107, 124]. We will unpack these ideas for the presented LZ model. The SLD operators $\hat{L}_{\omega}$ and $\hat{L}_{g}$ can be computed using Eq. (21), and take the concise analytical form:
@@
[…]
Since $\hat{L}_{\omega},\hat{L}_{g}]\neq 0$, there is no measurement basis that is optimal for both parameters simultaneously.This implies a fundamental quantum precision trade-off, as the information is encoded along non-orthogonal directions in Hilbert space that effectively “interfere” with each other. This interference prohibits a single measurement (or single-copy probe) from saturating the QCRB [122, 191, 192, 124]. Importantly, non-commutativity does not always imply incompatibility. In the LZ model we find that a particularly interesting condition holds:
@@ VII Quantum phase transitions, criticality, and critical metrology: many-body Ising model
The Landau–Zener model has served as a minimal framework for understanding the fundamental principles by which quantum critical systems acquire enhanced sensing capabilities. Having established these operating principles, we now turn to many-body quantum systems capable of supporting bona fide quantum phase transitions. For this purpose, we introduce the general form of a spin-chain Hamiltonian
$\hat{H}=\omega\sum_{i=1}^{N}\hat{\sigma}_{z}^{(i)}-\sum_{i\neq j}^{N}g_{ij}\hat{\sigma}_{x}^{(i)}\hat{\sigma}_{x}^{(j)}\;,$ (123)
where $\omega$ is the transverse field (the parameter to be estimated) and $g_{ij}$ denotes the interaction strength between spins $i$ and $j$. While Eq. (123) encompasses a broad class of interacting spin models, in what follows we restrict our analysis to two limiting cases. By first considering uniform nearest-neighbour interactions, $g_{ij}=g$, we obtain the one-dimensional transverse field Ising model (TFIM):
@@
  • Paramagnetic phase ($|\omega|>|g|$): In this phase, spins align along the transverse field direction $z$ with short-range correlations. Connected spin-spin correlations decay exponentially with distance, reflecting the absence of long-range order. Here, correlations between the spins are small, and the ground state is close to a product state. This is also known as the disordered phase.
  • Ferromagnetic phase ($|\omega|<|g|$): In this phase, spins tend to align along one of the two directions $\pm x$, reflecting the $\mathbb{Z}_{2}$ spin-flip symmetry of the Hamiltonian. For a finite system, the true ground state remains symmetry-preserving and is given by a superposition of the two ferromagnetic configurations. In the thermodynamic limit, however, spontaneous symmetry breaking occurs, and the system selects one of the ordered states. Connected correlations $\langle\hat{\sigma}_{x}^{(i)}\hat{\sigma}_{x}^{(i+r)}\rangle-\langle\hat{\sigma}_{x}^{(i)}\rangle\langle\hat{\sigma}_{x}^{(i+r)}\rangle$ saturate to a finite value as $r\to\infty$, signalling long-range order. This phase is also known as the ordered phase and supports a degenerate ground state manifold in the thermodynamic limit.
@@
The TFIM represents a class of systems which is exactly solvable via a Jordan-Wigner transformation by mapping spins to spinless fermions. A detailed discussion of the exact solution can be found in the following useful references [193, 194, 195, 176, 196, 197]. After Fourier transforming to momentum space, the Hamiltonian decouples into independent momentum-space pairs (k,k)(k,-k). These can be viewed as elementary quasiparticle modes that are always created or annihilated in pairs of opposite momentum. Within each such sector, the relevant states are the $|0\rangle$ (no quasiparticle excitations) and $\hat{c}^{\dagger}_{k}\hat{c}_{-k}^{\dagger}|0\rangle$ (quasiparticle pair with opposite momenta). This effectively reduces the problem to a two-level system for every momentum-$k$ mode. To make this structure explicit, we introduce the fermionic spinors
@@
In this language, the dynamics of each (k,k)(k,-k) block resembles a LZ type model whose pseudo-spin representation encodes the presence or absence of quasiparticle excitations. Here, the allowed momenta are given by:
@@
The $1/N$ scaling of the gap is a hallmark of the critical point in a finite-size system, reflecting how the system size itself sets the dominant length scale when the correlation length diverges. In this true many-body system, the closing of the gap emerges collectively from the full set of momentum-$k$ modes. In contrast to this, the LZ model provides a simplified analogue: its single avoided crossing mimics the role of these momentum-resolved level crossings, thereby capturing the essence of the true many-body critical behaviour. In this picture, each momentum block behaves like an effective two-level system whose ground state can be written as
@@
$|\psi_{GS}\rangle=\bigotimes_{k>0}|\psi_{GS}\rangle_{k}\;$ (132)
@@ VII.1 Criticality as a resource for sensing
We here seek to explore the underlying characteristics of critical systems which make them an appealing platform in the context of quantum sensing. Quantum criticality is driven by quantum fluctuations, as opposed to thermal fluctuations in classical phase transitions. At zero temperature changes in a Hamiltonian parameter, denoted $g$, causes a fundamental change in the ground state properties of the system. Notably, at or near the close vicinity of the critical point $g\equiv g_{c}$, the quantum fluctuations begin to dominate. This is accompanied by non-analytical behaviour in the ground state which is reflected in physical observables and qualitative changes in the nature of ground state correlations. At finite temperature, there is a delicate interplay between thermal and quantum fluctuations.
In such cases, there exists a finite “quantum critical” regime where quantum fluctuations still dominate before thermal effects become significant. This interplay has important consequences for the robustness and practicality of quantum critical sensors. Furthermore, quantum criticality and entanglement are closely connected [198, 199, 200, 201, 202], although no unified theory links them fully. At first glance, the genuine multipartite entanglement structure of ground states in quantum critical systems makes them attractive from a metrological perspective. For example, we have previously encountered the utility of leveraging entangled states in the context of metrology through Ramsey interferometry using entangled GHZ states in Section IV. In that instance, entanglement was an explicit resource facilitating enhanced sensing capabilities. However, as we will motivate in the following, it is the confluence of the entanglement structure in conjunction with the non-analytical behaviour near quantum critical points that underlies their utility as quantum sensors.
@@
in the traverse field $\omega$, manifest in large deviations of the ground state $|\psi_{GS}(\omega)\rangle$, with $\delta\omega$ representing a small perturbation. For this purpose, we will consider the fidelity of pure states: a geometric object quantifying the overlap between two quantum states $|\psi_{GS}(\omega)\rangle$ and $|\psi_{GS}(\omega^{\prime})\rangle$, defined as [203]
@@
The fidelity measures the closeness between the two quantum states. If the state is left unchanged then $F(\omega,\omega^{\prime})=1$, whereas for orthogonal states $F(\omega,\omega^{\prime})=0$. Contextually, the fidelity serves as a useful diagnostic tool for quantifying the variations $\delta\omega$ in the ground state properties across the transition point [28, 204, 26, 205, 206, 27, 207, 208].
Using the exact solution of the TFIM, the ground state fidelity takes the concise form [209]
@@
for arbitrary system size $N$, and where the product runs over momentum the blocks of the Hamiltonian $\hat{H}_{k}$. Fig. 12(a) shows the ground state fidelity (135) landscape as a function of the detuning from the critical point, $\delta g_{c}=g-g_{c}$, and system size $N$, for fixed $\omega=1$ and $\delta\omega=0.01$. The colour scale highlights two qualitatively different regimes: the extended yellow regions correspond to fidelity values very close to unity, indicating that the ground state remains largely unchanged under the small parameter shift. In contrast, the black/purple valley centred at the critical point, $\delta g_{c}=0$, marks a sharp loss in the fidelity (the ground state changes markedly) in the quantum-critical region. As the system size $N$ increases, this valley deepens and narrows, reflecting a growing instability of the ground state to perturbations at criticality. We further remark that in the thermodynamic limit, the fidelity can drop to zero irrespective of how small or large the deviation $\delta\omega$ is. This phenomena is known as the Anderson orthogonality catastrophe [210].
@@
Figure 12: Criticality as a metrological resource in the TFIM. (a) Ground state fidelity $F(\omega,\omega+\delta\omega)$ for $\delta\omega=0.01$ versus system size $N$ and detuning from critical point $\delta g_{c}=g_{c}-1$, with $\omega=g_{c}=1$. The sharp dip at $\delta g_{c}=0$ reflects maximal state distinguishability. Note the log-scale on the $y$-axis. (b) Fidelity at criticality $g_{c}=\omega=1$ decaying rapidly with system size NN (log-scale), signalling the ground state dramatical changing. (c) Fidelity susceptibility $\chi_{F}(\omega)$ (color in log-scale) as a function of system size NN (log-scale) and detuning from criticality $\delta g_{c}$. $\chi_{F}(\omega)$ diverges at the critical point marking the fact that the ground state changes dramatically. (d) Scaling of the fidelity susceptibility $\chi_{F}(\omega)$ at the critical point $g_{c}=\omega=1$. It shows a power-law scaling $\chi_{F}(\omega=g_{c})\propto N^{\mu}$, illustrating the diverging nature at the critical point.
@@ VII.2 Fidelity susceptibility
The connection between the dramatic loss of fidelity at the quantum critical point and the indication that these features are useful for metrology can be made rigorous by considering ground state properties which are differentiable in parameter space. In particular, the fidelity between two infinitesimally close ground states, for which $\delta\omega\ll 1$, can be expanded as [204, 28, 157, 207, 211, 209, 212]
@@
$\chi_{F}(\omega)$ $=-\frac{\partial^{2}F}{\partial(\delta\omega)^{2}}\biggr\rvert_{\delta\omega=0}$ (137)
@@
The fidelity $F(\omega,\omega^{\prime})$ is a geometric measure of how different the ground states are in parameter space, whereas the fidelity susceptibility $\chi_{F}(\omega)$ represents how sensitive the ground state is to infinitesimal variations in $\omega$. The fidelity susceptibility behaves like a response function in linear response theory. Concretely, it can be interpreted as a generalized susceptibility that quantifies the ground state’s response to an infinitesimal change in the parameter $\omega$ through its state overlap, rather than through changes in expectation values [213, 214, 205, 209, 28].
This is important as in quantum critical theory, linear response formalism is the standard tool to study how observables react to weak perturbations, with susceptibilities encoding correlations and critical scaling. In this sense, the fidelity susceptibility plays the role of a geometric analogue, quantifying the response of the ground state itself to infinitesimal parameter changes. For TFIM, the fidelity susceptibility can be computed directly
@@
We illustrate the behaviour of the fidelity susceptibility in Fig. 12(c), where it is shown as a function of the detuning from criticality, $\delta g_{c}=g-g_{c}$, and the system size $N$. In this analysis, the transverse field is fixed at $\omega=1$. The color-scale spans several orders of magnitude on a logarithmic scale: from dark purple (low susceptibility) of order $10^{-2}$ to bright yellow (high susceptibility) of order $10^{7}$, marking the large changes in the fidelity susceptibility. Several key features are immediately visible:
@@
  • Broadening away from criticality: For finite detuning ($|\delta g_{c}|>0$), the susceptibility grows much slower, indicating that the system is far less sensitive to parameter variations away from criticality.
@@
We can quantify the divergence of the fidelity susceptibility at the critical point by performing a scaling analysis. Setting $g=1$ for simplicity, we find that
$\chi_{F}$ $=\frac{1}{4}\sum_{k}\frac{\sin^{2}(k)}{\left[1-\cos(k)\right]^{2}}$ (141)
$\eqsim\frac{N}{2}\int_{\pi/N}^{\pi(N-1)/N}\frac{1}{k^{2}}dk$ (142)
@@
From this it is evident that $\chi_{F}(\omega)$ scales quadratically with system size $N$. Importantly, the fidelity susceptibility is not merely a heuristic measure of the ground state sensitivity to infinitesimal variations, but it can also be related to universal properties of a quantum critical point. Its scaling encodes universal properties of the system such as the correlation length exponent $\nu$—which governs the divergence of the correlation length, $\xi\sim|g-g_{c}|^{-\nu}$—and spatial dimension $d$ when evaluated at the transition point. In particular it has been shown [218, 207, 216, 215] that close to a quantum critical point, the fidelity susceptibility follows a universal power-law behaviour
@@
where the exponents depend only on the universality class of the transition, and not on the microscopic details of the Hamiltonian. The connection between the universal scaling functions of critical systems and the fidelity susceptibility is of direct consequence in metrology. This is because the fidelity susceptibility can in fact be seen as a specific component of the quantum geometric tensor, previously introduced in Section VI.4. Utilizing these connections, one finds that
@@
This makes the rigorous connection between the dramatic changes ground states undergo at quantum phase transitions, diverging susceptibilities, and ensuing metrological utility. Concretely, the fidelity susceptibility precisely captures the inherent change in ground state properties that emerge near the quantum critical point. Furthermore, $\chi_{F}$ directly encodes universal scaling features of a given phase transition. In turn, through the QCRB, the QFI describes the fundamental attainable precision in parameter estimation, and is directly related to the fidelity susceptibility. This therefore suggests that optimal measurement strategies utilizing critical quantum probes inherently leverage the diverging nature of critical systems. From this simple illustrative system, we have shown—through a geometric perspective—how features of critical systems naturally give rise to Heisenberg (quadratic) scaling in a metrological context. There exists an intrinsic connection between divergent susceptibilities and their role as metrological amplifiers in critical systems. It is important to note, however, that the arguments presented here apply strictly to the zero-temperature case. In more general settings, one expects a delicate interplay between thermal and quantum fluctuations, which jointly determine the metrological performance of quantum critical sensors, as illustrated in the LZ model discussed in Section VI.6.
The coveted Heisenberg scaling of precision in quantum measurements can be attained in both the Ramsey interferometric scheme utilizing entangled GHZ states (as we have seen in Section IV) and in the TFIM at the quantum critical point. Nevertheless, the underlying physical principles facilitating the attainability of this precision are fundamentally different. To clarify this distinction, one may regard the Ramsey protocol and the TFIM as paradigmatic examples of regular quantum metrology and critical quantum metrology, respectively. While the GHZ-based approach achieves enhanced sensitivity through maximally entangled probe sates and collective phase accumulation, critical quantum metrology explicitly leverages the diverging ground state properties and long-range correlations near quantum critical points as the driving metrological resource. The appealing aspect of critical quantum metrology is that such schemes are, for the most part, basically unencumbered by the practical necessities required to prepare highly entangled states explicitly and, furthermore, are more robust against decoherence effects which deteriorate any perceptible quantum advantage.